OpenStudy (anonymous):

If f(x) = x minus 3, all over 3, then what is f-1(x)

OpenStudy (anonymous):

answers: one-thirdx - 1 3x - 1 3x - 3 3x + 3

OpenStudy (campbell_st):

swap x and y so \[x = \frac{y - 3}{3}\] now make y the subject to get the answer for the inverse

OpenStudy (anonymous):

so 1/3 -1 ?

OpenStudy (campbell_st):

3x = y -3 y = 3x + 3

OpenStudy (anonymous):

oh i see thank you